# Patterns of active dipolar particles in external magnetic fields

### v2

@@ I Introduction
Systems of self-propelled particles as a prime example of active matter have been studied extensively for their complex collective behavior [1, 2, 3], both from the purely theoretical point of view where these systems provide examples for systems intrinsically out of equilibrium [4] and from an applied point of view where self-propelled particles are model systems for microrobots with biomedical and environmental applications [5]. From the applications point of view, the remote control of active particles is an important feature. An important method for remote control is through magnetic fields, provided that the active particles are equipped with a magnetic dipole moment [5, 6]. Active particles with a magnetic dipole moment can be found in nature, in the form of magnetotactic bacteria [6, 7], but can also be made fully synthetically, e.g. active magnetic colloids [8, 9] and various forms of magnetic propellers and microrobots [10, 11], and as biohybrid systems by magnetically functionalizing motile microorganisms [12, 13].
@@
While a large body of work has explored the magnetic steering of active particles as well as magnetically driven propulsion mechanisms [6], the effects of the magnetic interactions, both among the active particles and with an external magnetic field, have received considerably less attention. A few studies have explored the standard two-dimensional active Brownian particle model with additional magnetic interactions based on magnetic moments in the same plane as the active motion. Liao et al. [14] have determined the states of active particles with dipolar interactions in the absence of an external field and shown various patterns, in particular aggregation into chains. Similarly, our own earlier work has addressed patterns in small systems with only a few particles in confinement and studied the emergence of chain and rings [15]. These and additional structures were also found computationally as well as experimentally using centimeter-sized toy robots in a recent study with small particle numbers and harmonic confinement [16]. Moreover, motility-induced phase separation was shown to be suppressed by the magnetic interactions [14, 17]. Mixtures, in which dipolar particles are not active themselves, but interact with active particles, have also been studied and also show aggregation into chains, but also large-scale clustering of the particles [18, 19]. Finally, magnetic interactions were shown to result in the clustering of magnetotactic bacteria in flow through narrow channels [20, 21].
@@
In this study, we extend the previous work [14, 15] to study the patterns that emerge in systems of dipolar active particles in the presence of an external field. We use systematic computer simulations to study the ordered states in such system as a function of the density, field strength, dipole moment and activity. We distinguish up to eight different states based on three order parameters introduced by Liao et al. [14] that describe the competition between chaining, clustering and orientation. As a reference, we start with the case without the external field and study how a magnetic field modulates these states. We show that under strong fields, bands of oriented particles are formed that constitute active particle equivalents of columnar clusters know from passive ferrofluids [22, 23, 24].
@@ II.1 Model
$\xi^{\alpha}_{i}\!\left(t\right)\xi^{\beta}_{j}\!\left(t^{\prime}\right)\right>$ $=\delta_{i,j}\delta_{\alpha,\beta}\delta\!\left(t-t^{\prime}\right),$
@@
$\bm{F}_{i}=\sum_{i\neq j}\bm{F}^{\mathrm{dd}}_{ij}+\sum_{i\neq j}\bm{F}^{\mathrm{WCA}}_{ij}$
The pairwise forces between two particles $i$, $j$ separated by the distance $r_{ij}=\left|\bm{r}_{ij}\right|=\left|\bm{r}_{i}-\bm{r}_{j}\right|$ consist of dipolar forces and excluded volume interactions. The dipolar forces are given by [25]
@@
The excluded volume interactions are modeled by the Weeks-Chandler-Anderson (WCA) [26] pair potential with the interaction energy $\epsilon$ and the particle diameter $\sigma$
@@
The torques $\bm{\tau}_{i}$ acting on the particle $i$ result from the pairwise dipolar interactions [27] and the interaction with the external homogeneous magnetic field ${\bm{B}}=B\hat{\bm{e}}_{x}$ (stationary frame of reference),
$\bm{\tau}_{i}$ $=\sum_{i\neq j}\bm{\tau}^{\mathrm{dd}}_{ij}+\mu\hat{\bm{e}}_{i}\times{\bm{B}}$
$=\sum_{i\neq j}\frac{\mu_{0}\mu^{2}}{4\pi r_{ij}^{3}}\left[3\left(\hat{\bm{e}}_{i}\cdot\hat{\bm{r}}_{ij}\right)\left(\hat{\bm{e}}_{j}\times\hat{\bm{r}}_{ij}\right)+\left(\hat{\bm{e}}_{i}\times\hat{\bm{e}}_{j}\right)\right]+\mu\hat{\bm{e}}_{i}\times{\bm{B}}.$
@@ II.2 Simulation method
The dynamics of active dipolar particles were simulated using the simulation toolkit HOOMD-blue, version 2.6 [28, 29]. The active self-propulsion of the particles and their alignment in an external magnetic field were added through modules developed in-house. The project data and its parameter space was managed within a signac framework [30, 31].
All simulations used $N=1156$ particles and periodic boundary conditions. This relatively small system size was chosen to keep the simulation time manageable while systematically scanning a rather large parameter space. The same system size was used in previous work without an external field, where it was also shown that cluster sizes saturated for $N>1000$ [14]. To avoid interactions between periodic images of the particles, a cut-off distance of the order of half the system size, $r_{\mathrm{cut}}=L/2-\sigma$, was introduced for dipolar interactions. To recover correct rotational diffusion dynamics for spheres, we set $\gamma_{R}=3\gamma_{T}$ within HOOMD-blue. To study the effect of density, we varied the side length of the simulation box. We chose the following three densities. One well below ($\Phi=$0.13$,\,L=$85$$), one close to ($\Phi=$0.23$,\,L=$63$$) and one above ($\Phi=$0.57$,\,L=$40$$) the critical density where MIPS is expected in systems of active Brownian particles without dipolar interactions. To investigate the effect of a constant homogeneous external field on the collective dynamics of dipolar active particles, we systematically explored different combinations of the strength $B$ of the external field and the magnetic moment of the particles.
As in our previous work [15], we used the following dimensionless units: distances are measured in units of the particle size $\sigma$ and time in units of $\sigma^{2}/D_{0}^{T}$, the time in which a particle diffuses over a distance of its own size. The strength of the WCA interaction $\epsilon$ is used as the unit of energy, thus temperature $T$ has units of $\epsilon/k_{B}$ (here, $T=1$). All magnetic quantities are expressed relative to a reference field strength $B_{0}=1\times 10^{-5}$ T, which is of the order of the magnetic field of the Earth. Thus, the strength of the magnetic field has units of $B_{0}$ and magnetic moments are expressed units of $\epsilon/B_{0}$. Note that with these units, the self-propulsion velocity $v_{0}$ is identical to the Peclet number $v_{0}\sigma/D_{0}^{T}$. Numerical integration was performed for $N_{t}=10^{6}$ time steps with an integration time step width $\delta_{t}=2\times 10^{-5}$. Simulations were evaluated from $t=18$ to $t=20$ with data taken every $\Delta t=0.1$.
@@ II.3 Order parameters
In large systems of active dipolar particles, complex structures such as branches, connected networks, bands or loops within rings and large scale cluster formation are expected to emerge [32, 32, 14, 18, 19, 33]. To characterize such patterns, we used the following three global order parameters introduced by Liao et al. [14]. We used combinations of these three order parameters to identify and classify different state of the system and to construct state diagrams as described below.
@@
The fraction of the largest cluster is close to 0 when particles form small or no clusters and reaches 1 when the whole system forms one giant cluster. To identify clusters and, specifically, the largest cluster, we performed a density based cluster analysis (DBSCAN [34]) with the value of the cut-off distance $\delta_{c}$ based on the nearest neighbor distance $g_{1}(r)$ for each corresponding density.
@@
Finally, to assess the degree of chain formation in large systems, we applied the following chain criteria, following Liao et al. [14]. Two particles are considered to be bonded if they fulfill the following three criteria: First, their distance is smaller than the cut-off distance, $r_{ij}<\delta_{c}$. Second, they are aligned in parallel $\hat{\bm{e}}_{i}\cdot\hat{\bm{e}}_{j}>0$, and third their orientation is in line with their connecting distance vector $\left(\hat{\bm{e}}_{i}\times\hat{\bm{r}}_{ij}\right)\cdot\left(\hat{\bm{e}}_{j}\times\hat{\bm{r}}_{ij}\right)>0$. The fraction of particles that fulfill all three criteria give the degree of polymerization
@@ III Results
In this study, we investigated how a homogeneous external magnetic field affects structure formation and collective behavior of (magnetic) dipolar active particles. To this end, we performed extensive Brownian dynamics simulations of active dipolar particles in a homogeneous external magnetic field while systematically varying the field strength $B$, the active velocity $v_{0}$, and the magnetic moment $\mu$ (all in dimensionless units, see Methods). We performed our simulations for three representative densities $\Phi=$0.13 , 0.23 and 0.570.130.230.570.130.230.57start_ARG 0.13 end_ARG , start_ARG 0.23 end_ARG and start_ARG 0.57 end_ARG. We classified the different states of the system with respect to their clustering, orientational order, and chain formation, summarizing the results in state diagrams.
@@ III.1 No external field
As a reference, we first simulated active dipolar particles without an external magnetic field ($B=0$). Here, we briefly present our results, in particular how the observed states depend on the activity $v_{0}$ and the pairwise magnetic interactions via the magnetic moment $\mu$. This system has previously been analyzed in detail by Liao et al. [14] and the following results are in agreement with their results. In simulations for different $v_{0}$ and $\mu$ and different densities $\Phi$, we observe different patterns. Representative snapshots are shown in fig. 2. These snapshots reflect combinations of three different types of ordering: alignment of the magnetic moments, clustering of the particles, and chain formation. To quantify the order in these patterns, we make use of combinations of the three previously introduced global order parameters, the fraction of particles in the largest cluster $\left\langle\phi_{c}\right\rangle$, the global polarization $\left\langle\phi_{e}\right\rangle$, and the degree of polymerization $\left\langle\phi_{p}\right\rangle$ (see Methods). All three parameters are dimensionless quantities with values between zero and one. For a systematic classification, we consider the system as ordered if the corresponding order parameter exceed 0.5. This choice defines three criteria for order and combinations of these three criteria thus define eight possible states, as summarized in table 1.
@@
Figure 3: Diagram of states in the absence of an external field for three densities $\Phi=$0.13 , 0.23 and 0.570.130.230.570.130.230.57start_ARG 0.13 end_ARG , start_ARG 0.23 end_ARG and start_ARG 0.57 end_ARG (left to right). The states result from order parameter criteria combinations according to table 1. The red box in the right diagram ($\Phi=0.57$) highlights cluster formation that corresponds to MIPS for apolar ($\mu=0$) active particles.
@@
At the low and intermediate densities we simulated, we can distinguish six of these states that are shown in the snapshots in fig. 2: the disordered gas state with no order, the oriented gas with only orientational order, the gas of chains with only chain formation, oriented chains (chain formation and orientational order), a network of chains based on the combination of chain formation and clustering, and bands in the case where all three types of order are combined. At high system density, well above the critical density for motility induced phase separation (MIPS, $\Phi_{\textnormal{MIPS}}\approx 0.28$ [35]), we also see cluster formation in the apolar limit ($\mu=0$) that corresponds to MIPS, as expected [14]. The eighth possible state, oriented clusters, characterized by clustering and orientational order, but no chaining is not seen in the absence of an external field.
Based on the systematic classification, we summarized our results in the diagrams of states shown in fig. 3. For systems with densities below the critical density of MIPS ($\Phi=0.13$ and $\Phi=0.23$, fig. 3 left and center) and weak magnetic interactions ($\mu$ 1) no classification criterion is fulfilled and the system is in a disordered gas state (gray circle in fig. 3, snapshot in fig. 2 a). Order (chain formation and/or global polarization) emerges for stronger magnetic interactions ($\mu\geq 1.0$). At this point, contributions from dipolar interactions equal the energy contributed by thermal noise ($\mu^{2}=T=1$) and induce chain formation, consistent with previous observations in small systems [15]. Different ordered states can be realized by varying the activity $v_{0}$ for fixed magnetic moment. For example, in the case of intermediate magnetic interactions ($\mu=1.4$) systems with low levels of activity tend to form a gas of chains (pink triangle in fig. 3, snapshot in fig. 2 c). Increasing the activity adds global polarization to the chaining and, thus, results in oriented chains (green pentagon in fig. 3, snapshot in fig. 2 d). For even higher levels of activity, the formation of chains is suppressed while global polarization persists, so that an oriented gas is observed (purple plus in fig. 3, snapshot in fig. 2 b).
When the system density exceeds the critical density of MIPS ($\Phi=0.57$, fig. 3 right), apolar ($\mu=0$) and weakly magnetic particles ($\mu<1$) also show ordered states like a network of chains (yellow square in fig. 3, snapshot in fig. 2 e) or a gas of chains (pink triangle). In case of apolar particles, increasing the activity results in cluster formation (brown triangle) which has previously been identified as MIPS [14] (and was confirmed here by a bimodal distributions of the local density). At high system density, network of chains (yellow square) and bands (red cross in fig. 3, snapshot in fig. 2 f) dominate the diagram of states.

### v4

@@ I Introduction
Systems of self-propelled particles as a prime example of active matter have been studied extensively for their complex collective behavior [Vicsek2012, Romanczuk2012, Bechinger2016], both from the purely theoretical point of view where these systems provide examples for systems intrinsically out of equilibrium [Cates2015a] and from an applied point of view where self-propelled particles are model systems for microrobots with biomedical and environmental applications [Bente2018]. From the applications point of view, the remote control of active particles is an important feature. An important method for remote control is through magnetic fields, provided that the active particles are equipped with a magnetic dipole moment [Bente2018, Klumpp2019]. Active particles with a magnetic dipole moment can be found in nature, in the form of magnetotactic bacteria [Klumpp2019, Bazylinski2004], but can also be made fully synthetically, e.g. active magnetic colloids [baraban2012catalytic, alsaadawi2021control] and various forms of magnetic propellers and microrobots [ghosh2009controlled, vach2015fast], and as biohybrid systems by magnetically functionalizing motile microorganisms [carlsen2014magnetic, magdanz2013development].
@@
While a large body of work has explored the magnetic steering of active particles as well as magnetically driven propulsion mechanisms [Klumpp2019], the effects of the magnetic interactions, both among the active particles and with an external magnetic field, have received considerably less attention. A few studies have explored the standard two-dimensional active Brownian particle model with additional magnetic interactions based on magnetic moments in the same plane as the active motion. Liao et al. [Liao2020] have determined the states of active particles with dipolar interactions in the absence of an external field and shown various patterns, in particular aggregation into chains. Similarly, our own earlier work has addressed patterns in small systems with only a few particles in confinement and studied the emergence of chain and rings [Telezki2020]. These and additional structures were also found computationally as well as experimentally using centimeter-sized toy robots in a recent study with small particle numbers and harmonic confinement [obreque2024dynamics]. Moreover, motility-induced phase separation was shown to be suppressed by the magnetic interactions [Liao2020, sese2022impact]. Mixtures, in which dipolar particles are not active themselves, but interact with active particles, have also been studied and also show aggregation into chains, but also large-scale clustering of the particles [Maloney2020, Maloney2020a]. Finally, magnetic interactions were shown to result in the clustering of magnetotactic bacteria in flow through narrow channels [waisbord2016destabilization, meng2018clustering].
@@
In this study, we extend the previous work [Liao2020, Telezki2020] to study the patterns that emerge in systems of dipolar active particles in the presence of an external field. We use systematic computer simulations to study the ordered states in such system as a function of the density, field strength, dipole moment and activity. We distinguish up to eight different states based on three order parameters introduced by Liao et al. [Liao2020] that describe the competition between chaining, clustering and orientation. As a reference, we start with the case without the external field and study how a magnetic field modulates these states. We show that under strong fields, bands of oriented particles are formed that constitute active particle equivalents of columnar clusters know from passive ferrofluids [Liu1995, Flores1999, Ivey2001].
@@ II.1 Model
$\left<\xi^{\alpha}_{i}\!\left(t\right)\xi^{\beta}_{j}\!\left(t^{\prime}\right)\right$> $=\delta_{i,j}\delta_{\alpha,\beta}\delta\!\left(t-t^{\prime}\right),$
@@
$\bm{F}_{i}=\sum_{j\neq i}\bm{F}^{\mathrm{dd}}_{ij}+\sum_{j\neq i}\bm{F}^{\mathrm{WCA}}_{ij}$
The pairwise forces between two particles $i$, $j$ separated by the distance $r_{ij}=\left|\bm{r}_{ij}\right|=\left|\bm{r}_{i}-\bm{r}_{j}\right|$ consist of dipolar forces and excluded volume interactions. The dipolar forces are given by [Yung1998]
@@
The excluded volume interactions are modeled by the Weeks-Chandler-Anderson (WCA) [Chandler1983] pair potential with the interaction energy $\epsilon$ and the particle diameter $\sigma$
@@
The torques $\bm{\tau}_{i}$ acting on the particle $i$ result from the pairwise dipolar interactions [Yung1999] and the interaction with the external homogeneous magnetic field ${\bm{B}}=B\hat{\bm{e}}_{x}$ (stationary frame of reference),
$\bm{\tau}_{i}$ $=\sum_{j\neq i}\bm{\tau}^{\mathrm{dd}}_{ij}+\mu\hat{\bm{e}}_{i}\times{\bm{B}}$
$=\sum_{j\neq i}\frac{\mu_{0}\mu^{2}}{4\pi r_{ij}^{3}}\left[3\left(\hat{\bm{e}}_{i}\times\hat{\bm{r}}_{ij}\right)\left(\hat{\bm{e}}_{j}\cdot\hat{\bm{r}}_{ij}\right)-\left(\hat{\bm{e}}_{i}\times\hat{\bm{e}}_{j}\right)\right]+\mu\hat{\bm{e}}_{i}\times{\bm{B}}.$
@@ II.2 Simulation method
The dynamics of active dipolar particles were simulated using the simulation toolkit HOOMD-blue, version 2.6 [Anderson2008, Anderson2020]. The active self-propulsion of the particles and their alignment in an external magnetic field were added through modules developed in-house. The project data and its parameter space was managed within a signac framework [Adorf2018, Ramasubramani2018].
All simulations were carried out using $N=1156$ particles placed in a quadratic simulation box with side length $L_{x}=L_{y}=L$ with periodic boundary conditions. This relatively small system size was chosen to keep the simulation time manageable while systematically scanning a rather large parameter space. The same system size was used in previous work without an external field, where it was also shown that cluster sizes saturated for $N>1000$ [Liao2020]. To avoid interactions between periodic images of the particles, a cut-off distance of the order of half the system size, $r_{\mathrm{cut}}=L/2-\sigma$, was introduced for dipolar interactions. To recover correct rotational diffusion dynamics for spheres, we set $\gamma_{R}=3\gamma_{T}$ within HOOMD-blue. To study the effect of density, we varied the lateral size $L$ of the simulation box. We chose the following three density values defined by the mean area fraction $\Phi=N\pi\sigma^{2}/\left(4L^{2}\right)$: One well below ($\Phi=$0.13$,\,L=$85$\$), one close to ($\Phi=$0.23$,\,L=$63$\$) and one above ($\Phi=$0.57$,\,L=$40$\$) the critical density where MIPS is expected in systems of active Brownian particles without dipolar interactions. To investigate the effect of a constant homogeneous external field on the collective dynamics of dipolar active particles, we systematically explored different combinations of the strength $B$ of the external field and the magnetic moment of the particles.
As in our previous work [Telezki2020], we used the following dimensionless units: distances are measured in units of the particle size $\sigma$ and time in units of $\sigma^{2}/D_{0}^{T}$, the time in which a particle diffuses over a distance of its own size. The strength of the WCA interaction $\epsilon$ is used as the unit of energy, thus temperature $T$ has units of $\epsilon/k_{B}$ (here, $T=1$). All magnetic quantities are expressed relative to a reference field strength $B_{0}=1\times 10^{-5}$ T, which is of the order of the magnetic field of the Earth. Thus, the strength of the magnetic field has units of $B_{0}$ and magnetic moments are expressed units of $\epsilon/B_{0}$. Note that with these units, the self-propulsion velocity $v_{0}$ is identical to the Peclet number $v_{0}\sigma/D_{0}^{T}$. Numerical integration was performed for $N_{t}=10^{6}$ time steps with an integration time step width $\delta_{t}=2\times 10^{-5}$. Simulations were evaluated from $t=18$ to $t=20$ with data taken every $\Delta t=0.1$.
@@ II.3 Order parameters
In large systems of active dipolar particles, complex structures such as branches, connected networks, bands or loops within rings and large scale cluster formation are expected to emerge [Klapp2005, Klapp2005, Liao2020, Maloney2020, Maloney2020a, Holm2005]. To characterize such patterns, we used the following three global order parameters introduced by Liao et al. [Liao2020]. We used combinations of these three order parameters to identify and classify different state of the system and to construct state diagrams as described below.
@@
The fraction of the largest cluster is close to 0 when particles form small or no clusters and reaches 1 when the whole system forms one giant cluster. To identify clusters and, specifically, the largest cluster, we performed a density based cluster analysis (DBSCAN [Pedregosa2011]) with the value of the cut-off distance $\delta_{c}$ based on the nearest neighbor distance $g_{1}(r)$ for each corresponding density.
@@
Finally, to assess the degree of chain formation in large systems, we applied the following chain criteria, following Liao et al. [Liao2020]. Two particles are considered to be bonded if they fulfill the following three criteria: First, their distance is smaller than the cut-off distance, $r_{ij}<\delta_{c}$. Second, they are aligned in parallel $\hat{\bm{e}}_{i}\cdot\hat{\bm{e}}_{j}>0$, and third their orientation is in line with their connecting distance vector $\left(\hat{\bm{e}}_{i}\times\hat{\bm{r}}_{ij}\right)\cdot\left(\hat{\bm{e}}_{j}\times\hat{\bm{r}}_{ij}\right)>0$. The fraction of particles that fulfill all three criteria give the degree of polymerization
@@ III Results
In this study, we investigated how a homogeneous external magnetic field affects structure formation and collective behavior of (magnetic) dipolar active particles. To this end, we performed extensive Brownian dynamics simulations of active dipolar particles in a homogeneous external magnetic field while systematically varying the field strength $B$, the active velocity $v_{0}$, and the magnetic moment $\mu$ (all in dimensionless units, see Methods). We performed our simulations for three representative densities $\Phi=$0.13 , 0.23 and 0.570.130.230.57. We classified the different states of the system with respect to their clustering, orientational order, and chain formation, summarizing the results in state diagrams.
@@ III.1 No external field
As a reference, we first simulated active dipolar particles without an external magnetic field ($B=0$). Here, we briefly present our results, in particular how the observed states depend on the activity $v_{0}$ and the pairwise magnetic interactions via the magnetic moment $\mu$. This system has previously been analyzed in detail by Liao et al. [Liao2020] and the following results are in agreement with their results. In simulations for different $v_{0}$ and $\mu$ and different densities $\Phi$, we observe different patterns. Representative snapshots are shown in fig. 2. These snapshots reflect combinations of three different types of ordering: alignment of the magnetic moments, clustering of the particles, and chain formation. To quantify the order in these patterns, we make use of combinations of the three previously introduced global order parameters, the fraction of particles in the largest cluster $\left\langle\phi_{c}\right\rangle$, the global polarization $\left\langle\phi_{e}\right\rangle$, and the degree of polymerization $\left\langle\phi_{p}\right\rangle$ (see Methods). All three parameters are dimensionless, ranging from 0 to 1. For systematic classification we define a system as ordered when its order parameter exceeds 0.5, i.e., when a majority of particles adopt the corresponding ordering. This threshold provides a rigorous criterion for order. By combining the three binary (ordered/disordered) states, eight distinct state classifications arise, as summarized in table 1. We note that this classification is more detailed than the one used in previous work [Liao2020], even though it is based on the same order parameters. It allows us to systematically distinguish different types of order, which will become crucial in the presence of an external field as studied below.
@@
Figure 3: Diagram of states in the absence of an external field for three densities $\Phi=$0.13 , 0.23 and 0.570.130.230.57 (left to right). The states result from order parameter criteria combinations according to table 1. The red box in the right diagram ($\Phi=0.57$) highlights cluster formation that corresponds to MIPS for apolar ($\mu=0$) active particles.
@@
At the low and intermediate densities we simulated, we can distinguish six of these states that are shown in the snapshots in fig. 2: the disordered gas state with no order, the oriented gas with only orientational order, the gas of chains with only chain formation, oriented chains (chain formation and orientational order), a network of chains based on the combination of chain formation and clustering, and bands in the case where all three types of order are combined. At high system density, well above the critical density for motility induced phase separation (MIPS, $\Phi_{\textnormal{MIPS}}\approx 0.28$ [Bruss2017]), we also see cluster formation in the apolar limit ($\mu=0$) that corresponds to MIPS, as expected [Liao2020]. The eighth possible state, oriented clusters, characterized by clustering and orientational order, but no chaining is not seen in the absence of an external field.
Based on the systematic classification, we summarized our results in the diagrams of states shown in fig. 3. For systems with densities below the critical density of MIPS ($\Phi=0.13$ and $\Phi=0.23$, fig. 3 left and center) and weak magnetic interactions ($\mu<1$) no classification criterion is fulfilled and the system is in a disordered gas state (gray circle in fig. 3, snapshot in fig. 2 a). Order (chain formation and/or global polarization) emerges for stronger magnetic interactions ($\mu\geq 1.0$). At this point, contributions from dipolar interactions equal the energy contributed by thermal noise ($\mu^{2}=T=1$) and induce chain formation, consistent with previous observations in small systems [Telezki2020]. Different ordered states can be realized by varying the activity $v_{0}$ for fixed magnetic moment. For example, in the case of intermediate magnetic interactions ($\mu=1.4$) systems with low levels of activity tend to form a gas of chains (pink triangle in fig. 3, snapshot in fig. 2 c). Increasing the activity adds global polarization to the chaining and, thus, results in oriented chains (green pentagon in fig. 3, snapshot in fig. 2 d). For even higher levels of activity, the formation of chains is suppressed while global polarization persists, so that an oriented gas is observed (purple plus in fig. 3, snapshot in fig. 2 b).
When the system density exceeds the critical density of MIPS ($\Phi=0.57$, fig. 3 right), apolar ($\mu=0$) and weakly magnetic particles ($\mu<1$) also show ordered states like a network of chains (yellow square in fig. 3, snapshot in fig. 2 e) or a gas of chains (pink triangle). In case of apolar particles, increasing the activity results in cluster formation (brown triangle) which has previously been identified as MIPS [Liao2020] (and was confirmed here by a bimodal distributions of the local density). At high system density, network of chains (yellow square) and bands (red cross in fig. 3, snapshot in fig. 2 f) dominate the diagram of states.